The following table contains the number of successes and failures for three categories of a variable. Test whether the proportions are equal for each category at the α=0.01

level of significance.

Category 1

Category

Category 3

Failures

69

49

61

Successes

67

74

71

What is the​ P-value?_________​(Round to three decimal places as​ needed.)

please circle the answer
Answer #1

Ho: p1 = p2 = p3

H1: at least two proportions are unequal

Chi-Square Test of independence
Observed Frequencies
0
0 Category 1 Category Category 3 Total
Failures 69 49 61 179
Successes 67 74 71 212
Total 136 123 132 391
Expected frequency of a cell = sum of row*sum of column / total sum
Expected Frequencies
Category 1 Category Category 3 Total
Failures 136*179/391=62.261 123*179/391=56.309 132*179/391=60.43 179
Successes 136*212/391=73.739 123*212/391=66.691 132*212/391=71.57 212
Total 136 123 132 391
(fo-fe)^2/fe
Failures 0.7294 0.9488 0.0054
Successes 0.6159 0.8011 0.0045

Chi-Square Test Statistic,χ² = Σ(fo-fe)^2/fe =   3.105      
          
Level of Significance =   0.01      
Number of Rows =   2      
Number of Columns =   3      
Degrees of Freedom=(#row - 1)(#column -1) = (2- 1 ) * ( 3- 1 ) =   2      
          
p-Value =   0.212 [Excel function: =CHISQ.DIST.RT(χ²,df) ]  
Decision:    p value > α , do not reject Ho